OpenStudy (anonymous):

what is thesequence for 40 2010 5 ?

OpenStudy (anonymous):

40, 20, 10, 5

OpenStudy (anonymous):

see any pattern ?

OpenStudy (anonymous):

the question says find the eighth term of each sequence 40,20,10,5,......,......,......,....,....

OpenStudy (anonymous):

okay, do u see any pattern yet ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

take the ratios 20/40 = ? 10/20 = ? 5/10 = ?

OpenStudy (anonymous):

you wud get 1/2 for all above ratios

OpenStudy (anonymous):

that means its a geometric sequence with common ratio, r = 1/2 can you write the general term ?